OpenStudy (ammarah):

Write the standard form of the equation of the parabola with the given focus and vertex at (0,0).

OpenStudy (ammarah):

@agent0smith

OpenStudy (ammarah):

(2,0)

OpenStudy (ammarah):

@whpalmer4

OpenStudy (cwrw238):

if my memory serves me correctly the standard equation is y^2 = 2ax where the focus is at (a,0) so that would make it y^2 = 8x

OpenStudy (cwrw238):

no y^2 = 4ax!! giving y^2 = 8x

OpenStudy (cwrw238):

I'll check that out..

OpenStudy (cwrw238):